How many grams of hydrogen gas are required to react completely with oxygen gas and produce 24.5 G of water? 2H2 +O2 —->> 2H2O
1
Expert's answer
2020-04-22T10:54:49-0400

Molar mass of water =18grams

moles of water produced = "\frac{24.5}{18}" =1.36 moles

since from the equation we can see that stochiometric coefficient of hydrogen and water is in ratio 1:1

hence mole of hydrogen required will be equal to mole of water produced

mole of hydrogen required =1.36

mass of hydrogen required =2*1.36 =2.72 grams